OpenStudy (anonymous):

what is the estimate and the exact answer to: 7 3/8 x 24 =

OpenStudy (anonymous):

\[73/8\times24\] is that it?

OpenStudy (anonymous):

no its 7 3/8 times 24

OpenStudy (anonymous):

\[7^{3/8}\times24\]

OpenStudy (anonymous):

estimated exact value will be 49.787808

OpenStudy (the_fizicx99):

I think she's referring to a fraction? \[7 \frac{ 3 }{ 8 } = \frac{ (7 \times 8) + 3 }{ 8 }\] the denominator stays the same. \[\frac{ (7 \times 8) + 3 }{ 8 } = \frac{ 59 }{ 8 }\] So now you basically have \(\frac{59}{8}\) x \(\frac{24}{1}\) Remember that: \(\frac{a}{b} •\frac{c}{d} = \frac{ac}{bd}\)